OpenStudy (anonymous):

ok i need someone to help me. I have the answer but dont know how to make the square root nto a decimal. I will try and draw it

OpenStudy (anonymous):

a 24ft ladder is placed against a verticle wall. with the bottom of the ladder standing level 2ft fro the base of the building. how high up does the ladder reach. I got \[\sqrt{1}\] bt it has to b a decimal.how do I do that?

OpenStudy (vishweshshrimali5):



OpenStudy (vishweshshrimali5):

@tena32 I think your answer is wrong. I suggest you to give it another try.

OpenStudy (vishweshshrimali5):

Try to use pythogoras theorem, and if you did the same earlier then please check your calculations again.

OpenStudy (anonymous):

this site was running funny and it didn't type my other number in. it was supposed to be 23 not 2

OpenStudy (vishweshshrimali5):

Even then you have done some mistake. Let the required height be h. Then according to the figure (that I drew), since, the triangle is a right angled triangle, I have, By pythogoras theorem, (height)^2 + (base)^2 = (hypotenuse)^2 

OpenStudy (vishweshshrimali5):

@tena32, did you get it correct now ?

OpenStudy (usman1995):

u need to apply Pythagoras theorem...... uderoot(2^2+24^2) solve it u get 24.08 ft

OpenStudy (usman1995):

oh if it was supposed to be 23 instead of two just put 23 in Pythagoras theorem where i've put 2 and u'll get the right answer

OpenStudy (usman1995):

\[\sqrt{23^{2}+24^{2}}\]

OpenStudy (usman1995):

anwser is 33.24ft